If the contribution margin ratio is 0.60, targeted operating income is $100,000, and targeted sales volume in dollars is $500,000, then the degree of operating leverage is ________.

3.00 times

0.67 times

0.33 times

2.00 times
Answer #1

Contribution margin

= 500,000 * 0.60 = 300,000

Operating income = 100,000

Degree of operating leverage = Contribution margin/Operating income

= 300,000/100,000

= 3 times

Option A is the answer